### v2

# A Lesson in Splats: Teacher-Guided Diffusion for 3D Gaussian Splats Generation with 2D Supervision

Abstract
We introduce a diffusion model for Gaussian Splats, SplatDiffusion, to enable generation of three-dimensional structures from single images, addressing the ill-posed nature of lifting 2D inputs to 3D. Existing methods rely on deterministic, feed-forward predictions, which limit their ability to handle the inherent ambiguity of 3D inference from 2D data. Diffusion models have recently shown promise as powerful generative models for 3D data, including Gaussian splats; however, standard diffusion frameworks typically require the target signal and denoised signal to be in the same modality, which is challenging given the scarcity of 3D data. To overcome this, we propose a novel training strategy that decouples the denoised modality from the supervision modality. By using a deterministic model as a noisy teacher to create the noised signal and transitioning from single-step to multi-step denoising supervised by an image rendering loss, our approach significantly enhances performance compared to the deterministic teacher. Additionally, our method is flexible, as it can learn from various 3D Gaussian Splat (3DGS) teachers with minimal adaptation; we demonstrate this by surpassing the performance of two different deterministic models as teachers, highlighting the potential generalizability of our framework. Our approach further incorporates a guidance mechanism to aggregate information from multiple views, enhancing reconstruction quality when more than one view is available. Experimental results on object-level and scene-level datasets demonstrate the effectiveness of our framework.
Figure 1: (Left) Standard diffusion training is typically constrained to same-modality supervision. We introduce a framework that breaks this barrier by decoupling the sources of noised samples and supervision. Leveraging imperfect predictions of a pretrained feedforward 3D reconstruction module, we train a denoiser that operates in 3D, while being supervised by sparse 2D views. (Right) When paired with two different noisy teachers, our diffusion model enhances reconstruction quality across both objects and scenes, particularly in regions where deterministic models struggle due to reconstruction ambiguity.
1 Introduction
3D reconstruction from single 2D images is essential for computer vision applications, such as augmented reality, robotics, and autonomous vehicles, which rely on inferring 3D structures from limited viewpoints. Lifting 2D images to 3D is an ill-posed problem, because different 3D shapes can produce identical 2D projections.
Current approaches for 3D reconstruction from single images can be categorized into two main types: deterministic predictions and generative models, each with distinct limitations that stem from the ill-posed nature of inferring 3D structures from 2D images.
A prevalent approach in 3D reconstruction is to use deterministic neural networks to map input images to 3D representations, such as Neural Radiance Fields (NeRF) [33], 3D Gaussian Splats (3DGS) [39], and triplanes [68]. By leveraging differentiable rendering techniques, these methods can train directly from 2D images, circumventing the need for large volumes of 3D data, which are relatively scarce. This is advantageous because annotated 3D data is often difficult or impractical to obtain, especially for real-world applications. However, despite ongoing performance improvements, deterministic models remain inherently limited by the ambiguity in the 2D-to-3D mapping. These models cannot fully capture the range of possible 3D structures that correspond to a single 2D image, leading to overly smooth or blurred outputs when supervised by appearance-based losses for a specific target. This limitation constrains their ability to generate diverse and accurate 3D reconstructions, particularly when multiple plausible 3D structures could correspond to the same 2D input.
In contrast, generative models, especially diffusion models [15, 8], have recently shown strong potential in representing 3D data. Diffusion models are trained to progressively denoise corrupted versions of 3D data to generate 3D outputs that are likely under the training set distribution, either by directly operating in the 3D space [1, 30, 34] or in a lower-dimensional latent space [40, 51, 39]. These models have proven powerful for generating realistic data across various modalities. However, diffusion models for 3D generation face a fundamental limitation due to their training process, in which the denoising model is trained on noisy samples using their clean counterparts as supervision. This requirement demands a substantial amount of 3D data, making these models difficult to scale to real-world applications where 3D data is limited. Consequently, while diffusion models are highly effective, their dependency on large 3D datasets restricts their applicability for 3D generation when starting from 2D image data.
Some attempts have been made to bypass these limitations by training 3D generative models using multi-view images [46]. These models aggregate information across multiple views, structuring predictions in 3D space. However, they still fall short in generation quality compared to deterministic methods and require multiple views for each target object, which is often impractical in real-world applications. Thus, although both deterministic and generative models have made strides in 3D reconstruction, the field lacks scalable, high-performance solutions that can infer 3D structures from single 2D images. Research into training 3D diffusion models from single-image inputs remains underexplored, highlighting an important gap that our work aims to address.
In this work, we propose a novel training strategy that fundamentally revises the principles of diffusion model training by decoupling the denoised modality from the supervision modality. To clarify the challenge, we note that although a denoiser can be trained in the 3D modality and supervised with 2D images, diffusion models require the noisy signal and the clean signal to remain in the same modality—here, in 3D. However, since our supervision comes from 2D images, the key question becomes: how to create a noisy 3D signal suitable for denoising?
Our solution leverages powerful deterministic 3D reconstruction methods as “noisy teachers”. While deterministic models can produce 3D representations that are imperfect, as manifested by blurry renderings, they nonetheless provide a useful starting point in 3D. This approach is advantageous because the goal in diffusion training is to apply noise to the signal. By introducing enough noise beyond a critical timestep $t^{*}$, the difference between the noisy 3D signal provided by the deterministic model and the true, clean 3D structure becomes minimal. This "sweet spot" in noise level, inspired by techniques like SDEdit [60], allows us to use the deterministic 3D output to create a suitable noisy input for diffusion training. At this stage, we can train a denoiser to predict the clean signal as is often done for guidance using Tweedie’s formula [10, 22, 38], and we can supervise this prediction using image-based rendering losses.
Figure 2: Our proposed framework for noisy-teacher-guided training of a 3D Gaussian Splat (3DGS) diffusion model. Using a pre-trained deterministic predictor network for 3DGS, which we refer to as the “noisy teacher” (left), in stage 1 (top) we lift sampled views to generate an imperfect 3DGS prediction, providing noisy samples and supervision for the diffusion denoiser in 3DGS with additional image supervision. In stage 2 (bottom), we decouple the noisy samples from supervision and instead use the noisy teacher to generate noisy samples at noise levels $t>t^{*}$, with a multi-step denoising strategy generating high-quality predictions to facilitate image-only supervision. Both stages incorporate cycle consistency regularization. See text for further details.
However, this alone is not sufficient because if the denoiser only learns from timesteps $t>t^{*}$, it is bound to producing blurry outputs. To overcome this, we introduce a second key innovation: a multi-step denoising strategy that replaces the traditional single-step denoising framework. Specifically, starting from a noise level $t>t^{*}$, our model performs iterative denoising, akin to its behavior during inference, progressively reducing noise over multiple steps until reaching the clean 3D structure. This multi-stage approach allows the model to generate sharp and detailed reconstructions, as each denoising step moves closer to the true 3D target. By supervising the final output with image-based rendering losses, we enable the model to learn to produce high-quality outputs while updates propagate across the entire denoising sequence. Consequently, our approach opens up the capability to directly train 3D diffusion models from 2D supervision.
Notably, this strategy is extremely flexible and can utilize various 3D teacher models as long as the diffusion prediction can be supervised by them. In our experiments, we demonstrate this flexibility using two types of deterministic models: the popular Splatter Image [48] and the more advanced and recent Flash3D [47]. With these models, we train on single object and scene data, respectively. In both cases, our method significantly improves the performance of the base teacher model by $0.5-0.85$ PSNR. Additionally, our diffusion model facilitates the incorporation of additional views through guidance, further boosting performance compared to standard optimization.
2 Related Work
2.1 3D Reconstruction from Sparse Views with Deterministic Models
Recent research has focused on generating 3D content from images using deterministic feed-forward models [17, 48, 63, 47]. Notably, these methods rely solely on posed 2D views for training, rather than requiring 3D data, making them scalable for in-the-wild training. While deterministic models are relatively simple to design and train, they struggle to capture the inherent variability of possible solutions in 3D reconstruction, often leading to blurry reconstructions in regions with large potential variability. In this work, we advocate for a generative 3D diffusion model to enable richer and more complex representations. We use deterministic models [47, 48] as a starting point to generate noisy samples, which are then used to train our diffusion model.
2.2 3D Generation with Diffusion Models
Diffusion models have shown impressive generative capabilities across various domains, leading to significant interest in applying them to 3D content generation.
Diffusion Models Trained Directly on 3D Data. One line of research focuses on designing diffusion models that directly operate in 3D space. These models have been developed for various 3D representations, including point clouds [51, 31, 65], meshes [1, 30], 3D Gaussian splats [34, 39], and neural fields [7, 35, 9, 43, 4]. While effective, these methods assume the availability of high-quality 3D datasets in the target representation, which are often scarce and lack the breadth of real-world diversity. This data scarcity limits the generalization and applicability of these models, particularly in in-the-wild scenarios.
Leveraging 2D Diffusion Models for 3D Content Creation. To address the scarcity of 3D data, recent works have explored leveraging 2D-trained diffusion models to create 3D content. A prominent technique in this line is Score Distillation Sampling (SDS), which “lifts” 2D score predictions to a shared 3D representation [36, 14, 20, 62, 53, 25, 32, 37]. However, a key challenge here is achieving view coherence, as 2D models only access the visible parts of an object, leading to potential issues such as the notorious Janus problem. To mitigate this, view-aware diffusion models, condition the generation of target views on one or more source views, incorporating relative camera transformations for enhanced coherence [54, 28, 45, 57, 58, 5, 11, 42, 24, 59, 29, 16].
3D Diffusion Models Supervised by 2D Images. Our work aligns with a relatively underexplored area focused on training diffusion models that operate in 3D space but are supervised only with 2D images. Traditionally, in diffusion models, the supervision signal is provided in the same modality as the noisy samples. Holodiffusion [19] introduced a method to train a 3D diffusion model for feature voxel grids using 2D supervision. To address the discrepancy between the noised samples and the noised target distribution, they apply an additional denoising pass, encouraging the model to learn both distributions simultaneously.
In contrast, our approach minimizes the distribution discrepancy between teacher-induced noised samples and (unavailable) target noise samples by focusing on large noise values and refining lower-noise predictions through a multi-step denoising process. Several approaches [2, 49, 46], denoise multi-view images using a denoiser structured to predict a 3D representation, which is then rendered into 2D views. However, these methods inherently rely on the bijectivity of multi-view and 3D representations, which only hold with a substantial number of images. Additionally, because the images are noised independently, they may not coherently represent the noisy 3D structure, potentially harming consistency. Our proposed method, in contrast, directly denoises within the 3D representation while using 2D views for supervision, addressing both data scarcity and view coherence by explicitly working in 3D space.
3 Method
Problem Formulation. We tackle the problem of training a 3D diffusion model, $D_{\theta}(s_{t},t,x_{\text{src}})$, that maps $N$ noisy 3D Gaussian Splats, $s_{t}\in\mathbb{R}^{N\times d}$ to their clean version $s_{0}$. Each of the Gaussian Splats is of dimension $d$, representing properties such as center, covariance, opacity, and color. The model is conditioned on a single image $x_{\text{src}}$ and uses $k\geq 1$ additional views of the same content for supervision, $\{x_{\text{tgt}}^{v}\}_{v=0}^{k-1}$, without access to 3D ground truth. We assume access to a pre-trained deterministic model $s_{0}^{\text{teacher}}=T_{\phi}(x)$, trained on the same sparse view data, that reconstructs 3D Gaussian Splats from a single image—or we can train such a model ourselves. Our method employs this trained model as a noisy teacher, generating noisy samples to train the diffusion model, which is supervised by the target image set $\{x_{\text{tgt}}^{v}\}_{v=0}^{k-1}$.
Overview. Our pipeline operates in two stages. First, we bootstrap the diffusion model by supervising it with the noisy teacher’s predictions (Section 3.2). We then proceed to fine-tune the diffusion model using multi-step denoising and rendering losses (Section 3.1). Both stages are further equipped with a cycle consistency regularization described in Section 3.3. Although the bootstrapping stage precedes fine-tuning in the pipeline, we present it second in this manuscript to facilitate a smoother explanation of our core contributions. The model pipeline is depicted in Fig. 2.
3.1 Decoupling Noised Samples from Supervision with Multi-Step Denoising
Our approach to overcoming the aforementioned uni-modality limitation of diffusion model training is to decouple the source for the noisy samples from the supervision. Specifically, in standard diffusion training, noise is added to the target ground truth sample, which is then fed to the denoiser for recovering the clean target. Here, we do not have access to true 3D target data; instead, we replace it with a 3D prediction from a pretrained deterministic model. As previously discussed this model is limited in its ability to generate the diverse plausible 3D structures often resulting in blurry and imprecise predictions, thus we consider it to be a “noisy teacher”. A key insight is that while the noisy teacher does not produce 3D Gaussian Splats (3DGS) that are sufficient as a standalone solution, they are useful as a starting point in our proposed framework. We further take inspiration from [60], which finds that with enough noise, the data distribution of two modalities can overlap. Based on this, we choose a timestep $t^{*}$ such that for $t\geq t^{*}$, the noisy samples generated by the noisy teacher are likely to align with those that would have resulted from a forward noising process applied to the true, unknown ground truth 3DGS. Denoting these samples as
$s_{t}=\sqrt{\alpha_{t}}\,s_{0}^{\text{teacher}}+\sqrt{1-\alpha_{t}}\,\epsilon,$ (1)
With the input image $x_{\text{src}}\sim p_{\text{data}}$ sampled from the image dataset and noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ — these notations are omitted for brevity throughout the manuscript. One might be tempted to train the denoiser using the standard training objective:
$\mathbb{E}_{x_{\text{src}},t>t^{*},\epsilon}\left[\|s_{0}^{\text{teacher}}-D_{\theta}(s_{t},t,x_{\text{src}})\|_{2}^{2}\right].$ (2)
However, a problem remains: the noise $\epsilon$ is the noise added to the noisy teacher, so predicting it would not help since it is not the noise from the unknown true target. Instead, we utilize the fact that the predicted 3DGS representation $s$ can be differentiably rendered in arbitrary view directions $v$. Denoting this rendering operation as $\mathcal{R}(s,v)$, we can modify the training scheme to:
$\mathbb{E}_{x_{\text{src}},v\sim\mathcal{U}[k],t>t^{*},\epsilon}\left[\|x_{\text{tgt}}^{v}-\mathcal{R}(D_{\theta}(s_{t},t,x_{\text{src}}),v)\|_{2}^{2}\right].$ (3)
Yet, an issue still exists. By limiting our sample range of timesteps, we do not sample small noise levels, and as a result, the model cannot recover the fine details essential for successful reconstruction. Sampling smaller timesteps is not ideal, as the model would then be trained on noisy samples from the incorrect distribution.
To address this, we revise the standard single-step denoising training and instead employ multi-step denoising, sequentially applying the model with the appropriate time-step conditioning until reaching the final clean 3D prediction, $\hat{s}_{0}=D_{\theta}(\hat{s}_{1},1,x_{\text{src}})\circ\dots\circ D_{\theta}(s_{t},t,x_{\text{src}})$. Rendered towards a target view, the loss becomes:
$\mathcal{L}_{\text{mlt-stp}}=\mathbb{E}_{x_{\text{src}},v\sim\mathcal{U}[k],t>t^{*},\epsilon}\left[\lambda_{t}\|x_{\text{tgt}}^{v}-\mathcal{R}(\hat{s}_{0},v)\|_{2}^{2}\right],$ (4)
where $\lambda_{t}$ assigns a weight per denoising step. This multi-step denoising process mirrors the inference process but allows the network parameters to update. By training the model in this way, the 3D denoiser learns to handle 3D data directly, while still being supervised using widely available 2D datasets. Please refer to the implementation details 4.2 for a discussion regarding the computational efficiency of this unrolled optimization.
3.2 Noisy Teacher Bootstrapping
Training a 3D diffusion model directly using the multi-step denoising paradigm described above is computationally expensive. This is primarily due to the increased memory costs of maintaining gradients over multiple denoising steps in 3D space, which limits batch sizes and reduces efficiency. To address this, we propose avoiding this training approach from scratch by first bootstrapping our model using the noisy teacher.
Specifically, we generate noisy samples $s_{t}$ from the noisy teacher, as shown in Equation 1, and supervise the generated 3DGS both directly in 3D:
$\ell_{\text{3DGS}}=\|s_{0}^{\text{teacher}}-D_{\theta}(s_{t},t,x_{\text{src}})\|^{2},$ (5)
and in 2D through the image rendered from the generated 3DGS:
$\ell_{\text{image}}=\|x_{\text{tgt}}^{v}-\mathcal{R}(D_{\theta}(s_{t},t,x_{\text{src}}),v)\|_{2}^{2}.$ (6)
These losses are combined to form our overall bootstrapping objective:
$\mathcal{L}_{\text{bootsrap}}=\mathbb{E}_{x_{\text{src}},v\sim\mathcal{U}[k],t\sim\mathcal{U}[T],\epsilon}[\ell_{\text{3DGS}}+\ell_{\text{image}}].$ (7)
While the 3D supervision signal from the noisy teacher is not perfect, it is already in the 3D domain, making it computationally efficient. This setup allows for standard single-step denoising training, which is faster and less memory-intensive, with additional robustness introduced by the image-based supervision. Training the diffusion model in this way brings it to a performance level comparable to the base teacher model, preparing it for the multi-step training stage, where it can be fine-tuned to significantly surpass the base model’s performance.
3.3 Cycle Consistency Regularization
Both the bootstrapping and fine-tuning phases with multi-step denoising utilize the image rendering loss. Inspired by cycle consistency losses in unpaired image-to-image translation [67], we propose to further regularize the model using the generated output $\hat{s}_{0}$ by utilizing the rendered image $\hat{x}_{\text{tgt}}=\mathcal{R}(\hat{s}_{0},v_{\text{tgt}})$ to drive a second Gaussian Splats prediction, denoted as $\tilde{s}_{0}$. We then render this second prediction back to the source view to define our cycle consistency loss term:
$\mathcal{L}_{\text{cyc}}=\|x_{\text{src}}-\mathcal{R}(\tilde{s}_{0},v_{\text{src}})\|_{2}^{2}.$ (8)
Intuitively, this loss aims to constrain the predicted rendered view not only to match the target image in terms of appearance similarity, but also to be reliable enough to drive the generation of the source view. This loss is applied in both training stages. In the bootstrapping phase, the second splat prediction $\tilde{s}_{0}$ is generated through the noisy teacher, maintaining efficiency by only requiring one additional network pass. As shown in our ablation study, this loss improves the performance of the bootstrapping phase. We note that this technique could, in principle, also be used to improve the base model used as the noisy teacher, although this is beyond the scope of this work.
In the multi-step fine-tuning phase, however, our model already outperforms the noisy teacher (even without the cycle consistency loss), so lifting the predicted image to 3D via the noisy teacher is not meaningful. Instead, we apply the multi-step denoising process directly.
4 Experiments
Method 1-view Cars 1-view Chairs
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
SRN [44] 22.25 0.88 0.129 22.89 0.89 0.104
CodeNeRF [18] 23.80 0.91 0.128 23.66 0.90 0.166
FE-NVS [13] 22.83 0.91 0.099 23.21 0.92 0.077
ViewsetDiff w/o $\mathcal{D}$ [46] 23.21 0.90 0.116 24.16 0.91 0.088
ViewsetDiff w $\mathcal{D}$ [46] 23.29 0.91 0.094 - - -
PixelNeRF [61] 23.17 0.89 0.146 23.72 0.90 0.128
VisionNeRF [27] 22.88 0.90 0.084 24.48 0.92 0.077
NeRFDiff [12] 23.95 0.92 0.092 24.80 0.93 0.070
Splatter Image [48] 24.00 0.92 0.078 24.43 0.93 0.067
SplatDiffusion 24.84 0.93 0.077 25.21 0.93 0.066
Table 1: ShapeNet-SRN: Single-View Reconstruction (test split). Our method achieves better quality on all metrics on the Car split and Chair dataset, while performing reconstruction in the 3D space.
4.1 Experimental Setups
Datasets. We conduct experiments using two datasets: the object-level ShapeNet-SRN [6, 44] and the scene-level RealEstate10k [66]. ShapeNet-SRN comprises synthetic objects across various categories. In line with Splatter Image [48] and PixelNeRF [61], we focus on the cars and chairs classes. The resolution for ShapeNet-SRN dataset is $128\times 128$, and the Splatter Image model is employed as the teacher for the ShapeNet experiments. RealEstate10k consists of real-world video data captured in both indoor and outdoor environments. Following Flash3D [47], we use a resolution of $256\times 384$ for training in our experiments. The Flash3D model serves as the teacher to guide our diffusion model at the bootstrapping stage.
Method Memory Usage (GB) Model Size (MB)
PixelNeRF [61] 3.05 113
VisionNeRF [27] 6.42 1390
Splatter Image [48] 1.71 646
Ours 1.15 295
Table 2: Memory Footprint and Model Size.
Evaluation Metrics. We adopt PSNR, SSIM [52] and LPIPS [64] as metrics for the evaluation of the image reconstruction and novel view synthesis.
Memory Usage and Model Size We report both GPU memory consumption and model size. As shown in Tab.2, our model exhibits a significantly smaller size compared to VisionNeRF and Splatter Image. While PixelNeRF has a smaller model size, our approach achieves lower GPU memory consumption on the ShapeNet-SRN dataset.
4.2 Implementation Details.
Multi-step Denoising. We train the model using 4 NVIDIA A6000 GPUs. The computational efficiency is demonstrated in Tab. 2. During the bootstrapping stage (stage 1), a batch size of $100$ per GPU is employed to train the diffusion model under the guidance of the teacher model. Following this, in stage 2, multi-step denoising is performed using a DDIM sampler with 10 inference steps. To manage the increased computational complexity during this phase, the batch size is reduced to $10$.
Misc. We use Adam [23] as our optimizer with $\beta_{1}=0.9,\beta_{2}=0.999$. We use a total noising steps of $100$, with a linear scheduling, starting from $0.0001$ to $0.2$. For the bootstrapping stage, we use the teacher model to provide both supervision and noised samples. Instead of predicting the noises added to the splatters, our diffusion model denoises the noised inputs to clean samples directly. We set $t^{*}$ to be 20. For the architecture of diffusion model, we use the U-Net implementation from diffusers  [^1] https://huggingface.co/docs/diffusers. For the consistency branch, we use a denoising step of 10.
5 frames 10 frames $\mathcal{U}[-30,30]$ frames
Model PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Syn-Sin [55] - - - - - - 22.30 0.740 -
SV-MPI [50] 27.10 0.870 - 24.40 0.812 - 23.52 0.785 -
BTS [56] - - - - - - 24.00 0.755 0.194
Splatter Image [48] 28.15 0.894 0.110 25.34 0.842 0.144 24.15 0.810 0.177
MINE [26] 28.45 0.897 0.111 25.89 0.850 0.150 24.75 0.820 0.179
Flash3D [47] 28.46 0.899 0.100 25.94 0.857 0.133 24.93 0.833 0.160
SplatDiffusion 29.12 0.932 0.087 26.54 0.887 0.122 25.40 0.873 0.135
Table 3: Novel View Synthesis. Our model shows superior performance on RealEstate10k on small, medium and large baseline ranges.
Setting Novel view synthesis Source view synthesis
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
(a.1) Splatter Image 24.1992 0.9213 0.0843 31.1158 0.9808 0.0269
(a.2) feedforwad (our architecture) 19.9947 0.8613 0.1588 23.2363 0.9165 0.0955
(b.1) stage I w only rendering loss 18.8201 0.8415 0.1862 20.9767 0.8815 0.1535
(b.2) stage I w diffusion & rendering loss 22.6078 0.9046 0.1083 28.2025 0.9690 0.0411
(b.3) stage II w diffusion & rendering loss 23.1323 0.9116 0.1061 29.4463 0.9750 0.0358
(b.4) stage II w only rendering loss 24.4936 0.9264 0.0945 31.9839 0.9850 0.0233
(c.1) stage II w timestep from $[0,T]$ 24.6998 0.9233 0.0907 33.1196 0.9820 0.0205
(c.2) stage II w timestep from $[t^{*},T]$ 24.8324 0.9300 0.0871 33.4658 0.9883 0.0176
(d.1) stage I w/o consistency 22.6078 0.9046 0.1083 28.2025 0.9690 0.0411
(d.2) stage I w consistency 23.7293 0.9181 0.0979 29.9227 0.9774 0.0254
(d.3) stage I, stage II w consistency 24.9137 0.9332 0.0847 33.7061 0.9886 0.0153
Table 4: Ablations Studies on Single view Reconstruction, evaluated on the validation set of ShapeNet-SRN Cars
4.3 Image conditioned reconstruction
ShapeNet-SRN. We benchmark our diffusion model on the ShapeNet-SRN dataset, as presented in Tab. 1. Using only a single input view, our model achieves PSNR improvements of $0.84$ and $0.78$ on the cars and chairs splits, respectively, compared to the Splatter Image baseline.
For qualitative evaluation, we compare our method with Splatter Image, which serves as our teacher model in Fig. 1 and in Fig. 3. As seen in the first row (Fig. 3 (a)), images generated by Splatter Image occasionally exhibit artifacts and distortions. In contrast, our model generally produces more fine-grained structures and higher-quality details. Furthermore, as shown in Fig.  reffig:qual (b), the Gaussians generated by our model are denser and exhibit regular shapes, whereas those produced by Splatter Image tend to be oversized and less uniform.
RealEstate10K. We evaluate our method against recent state-of-the-art approaches on the real-world RealEstate10K dataset. As shown in Tab. 3, our model outperforms the teacher network, Flash3D, across three different evaluation settings, achieving an average PSNR improvement of $0.5$. The visual comparisons in Fig. 1 and Fig. 3 further demonstrates the superiority of our method, consistently producing cleaner images while Flash3D struggles in unseen regions, resulting in blurry artifacts.
4.4 Additional View Guidance
Unlike deterministic feedforward models, diffusion models have the distinct advantage of incorporating guidance. In our approach, we condition the prediction of Gaussian Splats parameters on a single input view and can optionally leverage a second view as guidance during the denoising process, following the Universal Guidance framework [3]. Detailed explanations and formulations of the guidance mechanism are provided in the supplementary material.
Table 5 compares our view guidance method to a 2-view 3DGS optimization procedure, as outlined by [21], which is initialized using the base model. Our diffusion model demonstrates a $0.2$ PSNR improvement when incorporating image guidance, with an additional $0.2$ PSNR gain achieved through Gaussian Splats optimization, consistently outperforming the Splatter Image baseline. While here we demonstrate guidance in a two-view settings, the guidance mechanism can naturally be extended to multiview scenarios, which we leave for future exploration.
Method GS optim Guidance PSNR SSIM LPIPS
Splatter Image 24.75 0.93 0.06
25.24 0.94 0.06
Ours 25.18 0.93 0.06
25.36 0.94 0.06
25.55 0.95 0.05
Table 5: Additional-view guidance. Evaluated on a subset of the car split, our diffusion-based model better utilizes an additional view through guidance compared to 3DGS optimization.
Figure 3: Qualitative results. (a) Qualitative comparison on the ShapeNet-SRN dataset for the Car and Chair categories. Our model produces views that are more faithful to the source image and better maintain plausibility. (b) Comparison of Gaussian Splat outputs between Splatter Image and our diffusion model shows that our model generates more regular patterns that closely follow the object surface. (c) Scene-level qualitative comparison on the RealEstate10K dataset demonstrates that our method produces more realistic results, particularly in ambiguous areas, such as the 2D edge separating the bed and the floor.
4.5 Ablation
We conducted a series of ablation studies on the ShapeNet-SRN cars dataset to measure the effect of various architectural designs on both novel and source view synthesis. The results are summarized in Tab. 4.
Architectural Comparison. To assess whether our improvements stem from the diffusion framework or architectural changes, we trained a feedforward model using the same U-Net architecture as our diffusion model, with the timestep-related layers removed to create a deterministic version. This feedforward network directly predict gaussian parameter from the input image as Splatter Image. Due to a much smaller model size, it performs significantly worse than Splatter Image (Tab. 4.a), underscoring that the diffusion approach is essential to the enhanced results we observe.
Bootstrapping (stage 1). Bootstrapping is necessary for the initialization of our diffusion model. As shown in Tab.4(b.1), it produces unsatisfactory results to directly train diffusion model without the teacher model as guidance because of the indirect cross-modality supervision. With the teacher guidance, the diffusion model can produce better results (Tab.4 (b.2)), but still bounded by the teacher’s performance.
Multi-step denoising (stage 2). In Stage 2, we found that the teacher model limits the performance of our model if we continue to use it as guidance (Tab.4 (b.3)) Instead, we fine-tune the model only with the rendering loss, allowing the model to explore how to improve the rendering performance from the ground truth images.
Cycle consistency. By introducing a feedback loop in which the predicted target view images are rendered back to the source view and supervised with the ground truth input image, we achieve performance improvements in both stages, as demonstrated in Tab. 4(d).
5 Conclusion and Limitations
In this work, we introduced a novel framework for training 3D diffusion models without requiring large-scale 3D datasets. By leveraging deterministic predictors as noisy teachers and using sparse 2D views for supervision, our approach enables effective training of 3D diffusion models with significant performance improvements.
Limitations. Our framework is flexible and could extend to various 3D representations; however, the current implementation relies on pixel-aligned 3D GS, inheriting certain limitations. Specifically, the uneven Gaussian distribution—where Gaussians concentrate on visible views with insufficient coverage in occluded regions—can lead to oversmoothness in novel views. Future work could address this limitation by adapting our framework to support alternative 3D representations, further enhancing its robustness and generalizability.
Acknowledgment
Or Litany is a Taub fellow and is supported by the Azrieli Foundation Early Career Faculty Fellowship. This research was supported in part by an academic gift from Meta.
Supplementary Material
Figure 4: Visualization of the denoising process of our diffusion models, trained on the Car and Chair categories of ShapeNet-SRN dataset.
6 Additional results
6.1 Additional qualitative results
We present visual comparisons of our method to PixelNeRF [61] and VisionNeRF [27] on ShapeNet-SRN Cars and Chairs in Fig. 5. More qualitative results from RealEstate10K dataset is in Fig. 6.
Figure 5: Additional qualitative results. Qualitative comparisons on the ShapeNet-SRN dataset for additional viewpoints and objects from the Car and Chair categories. Our model produces views that are more faithful to the source image and better maintain plausibility, while maintaining the fast rendering of Splatter Image.
6.2 Additional ablations
Feedforward vs Diffusion Model. To evaluate whether the observed improvements originated from the diffusion framework or architectural modifications, we trained a feedforward model by removing the time-conditioning layers from the U-Net architecture while preserving its overall structure. For comparison, we predicted Gaussian parameters from a single input image following the splatter image. The feedforward model exhibited significantly worse performance, which we attribute to its reduced size, resulting in limited representational capacity. From Tab. 6, we conclude that the diffusion framework is more suitable for such generation tasks compared to deterministic models, producing better results even with a smaller model size.
Setting PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Splatter Image 24.1992 0.9213 0.0843
Feedforwad 19.9947 0.8613 0.1588
Table 6: Feedforward model vs Splatter Image.
Choices of losses. Through experiments (Tab. 7), we found that it produces terrible results to directly train a diffusion model using rendering loss (both in stage 1 and stage 2), because the supervision indirectly comes from the rendered image instead of the denoised spaltters, which makes it hard for the diffusion model to learn the accurate distribution.
Stage 1 Stage 2 PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
R 16.7284 0.7836 0.3733
R 18.8201 0.8415 0.1862
D 21.3050 0.8965 0.1182
R + D 22.6078 0.9046 0.1083
R + D R + D 23.1323 0.9116 0.1061
R + D R 24.4936 0.9264 0.0945
Table 7: Ablation of losses at two stages. ‘R’ and ‘D’ represent rendering loss and diffusion loss, respectively.
For stage 1 training, the performance improves using teacher model as guidance and it reports the best results using both rendering loss and diffusion loss.
For stage 2 training, if we continue to use the diffusion loss, the teacher model will limit the performance of our diffusion model. Therefore, we only use rendering loss at stage 2, allowing the model to explore how to minimize the rendering loss and improve the rendering performance.
Weighted loss at different timesteps. The difficulty of prediction at different timesteps varies. Therefore, during the stage 2 training, we assign different weights to the rendering loss obtained at different timesteps and accumulate them for back-propagation throughout the denoising steps. The ablation results are in Tab. 8.
Figure 6: Additional qualitative results. Qualitative comparisons on RealEstate10K dataset.
Setting PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/o weighted loss 22.8848 0.9116 0.1044
w weighted loss 24.4936 0.9264 0.0945
Table 8: Ablations of weighted loss at different timesteps
7 Data details
7.1 ShapeNet-SRN Cars and Chairs
We adhere to the standard protocol for the ShapeNet-SRN dataset. Specifically, we use the provided images, camera intrinsics, camera poses, and data splits provided by [44] with a resolution of $128\times 128$. Our method is trained using relative camera poses. For single-view reconstruction, view 64 serves as the conditioning view, while for additional-view guidance, views 88 is used as guidance view. All remaining available views are treated as target views, where we compute novel view synthesis metrics.
7.2 RealEstate10K
We obtain 65,384 videos and their corresponding camera pose trajectories from the provided youtube links. Using these camera poses, we perform sparse point cloud reconstruction with COLMAP [41]. For evaluation, we adopt the test split provided by MINE [26] and follow prior work by assessing PSNR on novel frames that are 5 and 10 frames ahead of the source frame. Additionally, we evaluate on a randomly sampled frame within an interval of $\pm$30 frames, using the same frames employed in MINE’s evaluation. For evaluation, we use a total of 3,205 frames. The results presented in Tab. 3 are sourced from Flash3D [47]. Our model is trained and tested at a resolution of $256\times 384$.
8 Implementation details
Multi-step Denoising. We train the model on 4 NVIDIA A6000 GPUs. Our diffusion model is quite efficient . For bootstrapping at stage 1, we use a batch size of $100$ on each GPU. After obtaining the diffusion model from the teacher model, we perform multi-step denoising with a DDIM sampler of 10 inference steps. The batch size for stage 2 reduces to 10. We assign different weights to the rendering loss obtained at different timesteps and accumulate them for back-propagation throughout the denoising steps.
Additional-view guidance Different from deterministic feedforward models, one significant advantage we gain from diffusion models is the ability of using guidance. We use one input view as the condition to predict the Gaussian Splats parameters and then use a second view as guidance during the denoising process using the forward guidance from Universal Guidance [3].
Since we predict $\hat{s}_{0}$ directly, the noise can be calculated as follows:
$\epsilon_{t}=\frac{s_{t}-\sqrt{\alpha_{t}}\hat{s}_{0}}{\sqrt{1-\alpha_{t}}}$ (9)
Then we calculate the gradient using the guidance image $x_{gd}$ and the corresponding view direction $v$:
$\text{grad}\leftarrow\nabla_{s_{t}}\ell\big{[}x_{gd},\mathcal{R}(\hat{s}_{0},v)\big{]}.$ (10)
With the guidance strength factor $s(t)$, we can obtain $\hat{\epsilon}_{t}$
$\hat{\epsilon}_{t}=\epsilon_{t}+s(t)\cdot\text{grad}$ (11)
At last, we can get $s_{t-1}$ following DDIM sampling:
$s_{t-1}=\sqrt{\alpha_{t-1}}\hat{s}_{0}+\sqrt{1-\alpha_{t-1}}\cdot\hat{\epsilon}_{t}$ (12)
Method GS optim Guidance PSNR SSIM LPIPS
Splatter Image 24.75 0.93 0.06
25.24 0.94 0.06
Ours 25.18 0.93 0.06
25.26 0.94 0.06
25.36 0.94 0.06
25.55 0.95 0.05
Table 9: Additional-view guidance. Evaluated on a subset of the car split, because per-sample GS optimization takes time.